# Kerr-MOG-(A)dS black hole and its shadow in scalar-tensor-vector gravity theory

### v2

@@ Abstract
The scalar-tensor-vector gravity (STVG) theory has attracted significant interest due to its ability to effectively address the issue of galaxy rotation curves and clusters of galaxies without considering the influence of dark matter. In this paper, we construct rotating black hole solutions with a cosmological constant in the STVG theory (i.e., Kerr-MOG-(A)dS black hole solutions), where the import of a gravitational charge as a source modifies the gravitational constant, determined by $G=G_{\text{N}}(1+\alpha)$. For Kerr-MOG-dS spacetime, the observer is situated at a specific location within the domain of outer communication, rather than being located infinitely far away. Since black hole shadows are shaped by light propagation in spacetime, the interaction between the MOG parameter and the cosmological constant is expected to produce novel effects on these shadows. As the cosmological constant $\Lambda$ increases, the apparent size of the black hole shadow on the screen decreases. Additionally, the shadow expands with an increase in the MOG parameter $\alpha$, reaching a maximum at a certain value, and its shape becomes more rounded under an arbitrary rotation parameter, which leads to degeneracy between different black hole parameters. However, by employing numerical backward ray-tracing techniques, we have found that gravitational lensing and the frame-dragging effect effectively distinguish this degeneracy. Our work contributes to a deeper understanding of black holes in modified gravity, their observational signatures, and constraints.
@@ I Introduction
On the observational side, despite numerous attempts to detect dark matter, all experimental efforts up to now have yet to achieve recognized success An et al. (2023); Ning et al. (2023). Therefore, on the theoretical side, STVG theory has attracted a great deal of attention because it not only successfully explains the rotation curves of galaxies and the properties of galaxy clusters, but also does so without considering dark matter Moffat and Rahvar (2013, 2014); Moffat and Toth (2015); Moffat (2014); Rahvar (2022). In addition, observational data from the Event Horizon Telescope (EHT) on the shadows of Sgr A* and M87* Akiyama et al. (2019a, b, 2022a) support the predictions of STVG theory and there are criteria to distinguish this theory from general relativity Moffat and Toth (2020). The axially symmetric solution in STVG theory provided by Moffat et al. is called the Kerr-MOG black hole Moffat (2015a). To date, several new black hole solutions related to this theory have been obtained Moffat (2021); Moffat and Toth (2021); Liu et al. (2023a). Building on the established solutions in STVG theory, extensive research has been conducted on various aspects of black holes, including their thermodynamic properties Mureika et al. (2016), geodesic motion Lee and Han (2017), shadows Moffat (2015b); Wang et al. (2019); Wei and Liu (2018); Guo et al. (2018); Sheoran et al. (2018); Zhang et al. (2024); Qiao et al. (2020); Konoplya and Zhidenko (2021); Younsi et al. (2016), and quasi-normal modes Manfredi et al. (2018); Jiang et al. (2024), etc.
@@
In this work, we aim to extend the Kerr-MOG solution Moffat (2015a) to more general cases with a cosmological constant, thereby obtaining Kerr-MOG-(A)dS black holes. Given the recent groundbreaking observations by the EHT, theoretical research on the properties of black holes frequently focuses on the analysis of black hole images Chen et al. (2023). Hence, we have also investigated the effects of the MOG parameter and a positive cosmological constant on black hole shadows, as well as the numerical estimation of the angular radius of the supermassive black holes Sgr A* and M87* within this theoretical framework. The remaining part of this paper is organized as follows. In Sec. II, we provide a concise overview of the STVG theory and give the Kerr-like solution of the field equations with a cosmological constant, i.e., Kerr-MOG-(A)dS black holes. In Sec. III, utilizing the geodesic equation, we derive the orbital equations for photons in the context of Kerr-MOG-dS spacetime. In Sec. IV, we examine the shadows of black holes, with a primary focus on the apparent shapes and distortions of the shadow. Also, we numerically calculate these two observables, $R_{s}$ and $\delta_{s}$, to characterize the shadow. Sec. V focuses on providing an overview of our conclusions while also contemplating the possible development and exploration of related research topics in the future.
@@ II.2 Rotating-(A)dS solutions
where $r_{m}$, $r_{h}$, and $r_{c}$ correspond to the Cauchy horizon, event horizon, and dS cosmological horizon, respectively Liu et al. (2023a). The parameters $r_{b}$, $r_{m}$, $r_{h}$, and $r_{c}$ satisfy the condition $r_{b}+r_{m}+r_{h}+r_{c}=0$. Moreover, $r_{c}\geq r_{h}>r_{m}>0$ and $r_{b}$ is negative.
@@ III Photon orbits
where $\tau$ is an affine parameter along the null geodesics, and $\mathcal{S}$ is the Jacobi action. We can set up the Killing fields $\xi^{a}=(\partial/\partial t)^{a}$ and $\psi^{a}=(\partial/\partial\varphi)^{a}$, then the corresponding energy $\mathcal{E}$ and the angular momentum in the direction of the axis of symmetry $L$ can be defined as Wald (1984); Zeng et al. (2022); Liu et al. (2023b):
@@ IV SHADOWS OF BLACK HOLES
While AdS solutions play a crucial role in AdS/CFT correspondences, they lack empirical support from specific astronomical observations. Therefore, in this section, we primarily focus on the black hole shadows in Kerr-MOG-dS spacetime. For this, we need to adopt the analysis approach of A. Grenzebach Grenzebach et al. (2014) and the numerical backward ray-tracing method Wang et al. (2018a, 2021); Hu et al. (2021); Zhong et al. (2021); Chen and Jing (2024) to investigate the shadow of Kerr-MOG black holes in the presence of a positive cosmological constant.
@@ IV.1 Apparent shape
Figure 1: Shadows cast by black holes with a cosmological constant, as seen by an observer at $r_{0}=50$ and $\vartheta_{0}=\pi/2$. The (a) and (c) are static cases $a/M=0$, corresponding to $\Lambda=10^{-4}$ and $\Lambda=10^{-3}$, respectively. The (b) and (d) are rotating case $\alpha=0.02$, corresponding to $\Lambda=10^{-4}$ and $\Lambda=10^{-3}$, respectively.
@@
Figure 2: Shadows cast by black holes with a spin $a/M=0.99$, as seen by an observer at $r_{0}=50$ and $\vartheta_{0}=\pi/2$. The (a), (b), (c) and (d) corresponding to $\alpha=0$, $\alpha=1$, $\alpha=2$ and $\alpha=3$, respectively.
@@
Figure 3: Shadows cast by black holes with a cosmological constant, as seen by an observer at $r_{0}=50$ and $\vartheta_{0}=\pi/2$. The (a) and (c) are spin $a/M=0.5$, corresponding to $\Lambda=10^{-4}$ and $\Lambda=10^{-3}$, respectively. The (b) and (d) are spin $a/M=0.99$, corresponding to $\Lambda=10^{-4}$ and $\Lambda=10^{-3}$, respectively.
@@
For simplicity, we consider fixing the position of the observer at $r_{0}=50$ and $\vartheta_{0}=\pi/2$. Where the angle $\vartheta_{0}$ is defined as the inclination angle between the observer’s line of sight and the rotation axis of the black hole. Form Eqs. (47) and (48) one can depict the contour of the black hole shadow under varying MOG parameter $\alpha$, spin $a$, or cosmological constant $\Lambda$. In Fig. 1, the shadows of static black holes and rotating black holes are shown under $\Lambda=10^{-4}$ or $\Lambda=10^{-3}$. When the cosmological constant is relatively small, the shadow scale exhibits an increase with the growth of the MOG parameter. However, as the cosmological constant increases, this trend changes. This shift can be attributed to the observer’s position $r_{0}$ approaching the cosmological horizon $r_{c}$, causing the observer’s position to have a more significant effect on the shadow size than the effects of the MOG modification. In Fig. 2, we present the shadow contour of a near-extreme black hole with different $\alpha$ and $\Lambda$, i.e., $a/M=0.99$. The results show that, as the MOG parameter increases, the range of the photon sphere radius narrows, and its contour increasingly resembles a standard circle. When the cosmological constant is relatively small, the shadow size increases with the growth of the MOG parameter. For a fixed $a$ and $\alpha$, the apparent shape of the shadow varies with $\Lambda$. As $\Lambda$ decreases, the shadow’s size expands. In Fig. 3, the same phenomenon that is clearly visible in Fig. 1-(c) can also be observed in rotating black holes. Furthermore, when the MOG parameter increases significantly, the rotation-induced distortion of the photon orbits becomes less noticeable. It is worth noting that in Figs. 1-3, the purple solid line represents the shadow of the Schwarzschild black hole, serving as a scale reference.  [^1] The coordinates $x/M$ and $y/M$ in Figs. 1-3 represent the on-screen scale, which is proportional to the actual celestial coordinate scales $\bar{x}/M$ and $\bar{y}/M$ as described by the geometric relationships in equations (47) and (48), depending on the observer’s position $r_{0}$. For example, in the Schwarzschild case, where $r_{0}=50$, the relation $\bar{x}\simeq 50.897x$ holds.
@@
The contours of static versus rotating black holes will be extremely similar when the MOG parameter is sufficiently large Camilloni et al. (2024), making it difficult to infer black hole parameters solely from the black hole shadow. However, does the frame-dragging effect caused by rotation in black hole spacetimes also affect the MOG parameter? For this, we can use the numerical backward ray-tracing method, which adopts the same orthonormal tetrad as the analytical methods in this paper and sets the number of pixels to $n=2048$. Figs. 4 and 5 effectively demonstrate the distortion of space by MOG black holes and the gravitational lensing effects these objects produce. Specifically, the gravitational lensing observed can be utilized to examine theories of gravity Liu et al. (2024); Chowdhuri et al. (2023); Ghosh and Bhattacharyya (2022). We can observe that although increasing the MOG parameter makes the contours of rotating black holes approach those of static black holes, it does not affect the frame-dragging effect induced by rotation. Moreover, Figs. 4 and 5 serve as double checks for Figs. 1-(a) and 2-(b), respectively, with each displaying the same black hole contours.
@@
Figure 4: Shadows cast by static black holes ($a/M=0$) with a cosmological constant, as seen by an observer at $r_{0}=50$ and $\vartheta_{0}=\pi/2$.
@@
Figure 5: Shadows cast by extremely rotating black holes ($a/M=0.99$) with a cosmological constant, as seen by an observer at $r_{0}=50$ and $\vartheta_{0}=\pi/2$.
@@ IV.2 Observable measurements
Up to this point, we have determined that parameters such as spin $a$, MOG parameter $\alpha$, and cosmological constant $\Lambda$ will significantly affect the apparent shape of the shadow in Kerr-MOG-dS spacetime. To characterize the shadow, it is essential to have easily measurable and reliable properties of observable measurements. We adopt two such observable measurements: the on-screen radius $R_{s}$ and distortion parameter $\delta_{s}$, following an approach analogous to the one taken in Ref. Hioki and Maeda (2009).
Figure 6: Shadows cast by black holes, as seen by an observer at $r_{0}=50$ and $\vartheta_{0}=\pi/2$. The observable measurements for the apparent shape of the black hole are the radius of the reference circle $R_{s}$ and the distortion parameter $\delta_{s}=\mathrm{D}/R_{s}$, where $\mathrm{D}$ is the difference between the left endpoints of the reference circle and of the shadow.
The parameter $R_{s}$ characterizes the on-screen size of the shadow, corresponding to the radius of the reference circle depicted as a blue dashed line in Fig. 6. The circle passes through three points: the top position $P_{A}=(x_{t},y_{t})$, the bottom position $P_{B}=(x_{b},y_{b})$, and the point $P_{C}=(x_{r},0)$, which corresponds to the unstable retrograde circular orbit as observed from the equatorial plane by an observer. Additionally, the point $P_{O}=(x_{o},0)$ represents the center of the reference circle, and $x_{o}$ can be obtained from the coordinates of $P_{A},P_{B}$, and $P_{C}$:
@@
Figure 7: observable measurements $R_{s}$ (left) and $\delta_{s}$ (right) as functions of $\alpha$ are presented, respectively, for $\Lambda=10^{-4}$ (top) and $\Lambda=10^{-3}$ (bottom). In all cases, the observer is placed at $r_{0}=50$ and $\vartheta_{0}=\pi/2.$
We numerically calculated these two observable measurements under the condition of fixing the observer position at $r_{0}=50$ with an inclination angle of $\vartheta_{0}=\pi/2$. The results are presented in Fig. 7. Note that the static case is not included in the results because the contour of the shadow coincides with the reference circle, resulting in no distortion. Figs. 7-(b) and 7-(d) show that the cosmological constant $\Lambda$ has little effect on $\delta_{s}$. But $\delta_{s}$ decreases as the parameter $\alpha$ increases, which suggests that the Kerr-MOG-dS black hole gets less deformed compared to the Kerr-dS black hole. Figs. 7-(a) and 7-(c) clearly show the radius $R_{s}$ of the shadow. From these results, it becomes evident that when the cosmological constant $\Lambda=10^{-4}$, $R_{s}$ increases as the parameter $\alpha$ increases with $\alpha\in[0,3]$. However, when the cosmological constant $\Lambda=10^{-3}$, there is a maximum value for $R_{s}$ , corresponding to $\alpha\approx 1.815$. This is consistent with the peculiar phenomena observed in Figs. 1-(c), 3-(c), and 3-(d) describing the shadow cast by MOG-dS black holes in Sec. IV.
For a approximatively estimation, utilizing the Kerr-MOG-dS metric (22), we calculate the angular radius of a black hole shadow, defined as $\theta_{\text{BH}}=\bar{R}_{s}\frac{\mathcal{M}}{D_{O}}$, with $D_{O}$ representing the distance from the observer to the black hole. The radius $\bar{R}_{s}$, which is proportional to the on-screen radius $R_{s}$ as described by the geometric relationships in equations (47) and (48) (see detail in footnote 1). Specifically, for a black hole possessing a mass of $\mathcal{M}$ and situated $D_{O}$ away from the observer, we express the angular radius $\theta_{\text{BH}}$ quantitatively as $\theta_{\text{BH}}=9.87098\times 10^{-6}\bar{R}_{s}\left(\frac{\mathcal{M}}{M_{\odot}}\right)\left(\frac{1\text{kpc}}{D_{O}}\right)\mu\text{as}$ Amarilla and Eiroa (2012).
@@
$\Lambda=0.0000$ 24.737 26.747 28.738 30.712 32.670
$\Lambda=0.0001$ 23.644 25.562 27.460 29.342 31.208
$\Lambda=0.0002$ 22.497 24.317 26.118 27.902 29.671
$\Lambda=0.0003$ 21.288 23.004 24.701 26.382 28.047
$\Lambda=0.0000$ 24.751 26.760 28.750 30.722 32.679
$\Lambda=0.0001$ 23.658 25.573 27.470 29.351 31.215
$\Lambda=0.0002$ 22.510 24.328 26.127 27.909 29.677
$\Lambda=0.0003$ 21.298 23.013 24.709 26.388 28.053
@@
$\Lambda=0.0000$ 19.859 21.473 23.081 24.656 26.229
$\Lambda=0.0001$ 18.982 20.522 22.046 23.556 25.055
$\Lambda=0.0002$ 18.062 19.522 20.968 22.401 23.821
$\Lambda=0.0003$ 17.090 18.457 19.816 21.164 22.517
$\Lambda=0.0000$ 19.871 21.484 23.064 24.664 26.236
$\Lambda=0.0001$ 18.993 20.531 22.054 23.563 25.061
$\Lambda=0.0002$ 18.071 19.531 20.975 22.406 23.826
$\Lambda=0.0003$ 17.099 18.475 19.837 21.185 22.521

### v3

@@ Abstract
The scalar-tensor-vector gravity (STVG) theory has attracted significant interest due to its ability to effectively address the issue of galaxy rotation curves and clusters of galaxies without considering the influence of dark matter. In this paper, we construct rotating black hole solutions with a cosmological constant in the STVG theory (i.e., Kerr-MOG-(A)dS black hole solutions), where the import of a gravitational charge as a source modifies the gravitational constant, determined by $G=G_{\text{N}}(1+\alpha)$. For Kerr-MOG-dS spacetime, the observer is situated at a specific location within the domain of outer communication, rather than being located infinitely far away. Since black hole shadows are shaped by light propagation in spacetime, the interaction between the MOG parameter and the cosmological constant is expected to produce novel effects on these shadows. As the cosmological constant $\Lambda$ increases, the apparent size of the black hole shadow decreases. Additionally, the shadow expands with an increase in the MOG parameter $\alpha$, reaching a maximum at a certain value, and its shape becomes more rounded under an arbitrary rotation parameter, which leads to degeneracy between different black hole parameters. However, by employing numerical ray-tracing techniques, we have found that gravitational lensing and the frame-dragging effect effectively distinguish this degeneracy. Our work contributes to a deeper understanding of black holes in modified gravity, their observational signatures, and constraints.
@@ I Introduction
On the observational side, despite numerous attempts to detect dark matter, all experimental efforts up to now have yet to achieve recognized success An et al. (2023); Ning et al. (2023). Therefore, on the theoretical side, STVG theory has attracted a great deal of attention because it not only successfully explains the rotation curves of galaxies and the properties of galaxy clusters, but also does so without considering dark matter Moffat and Rahvar (2013, 2014); Moffat and Toth (2015); Moffat (2014); Rahvar (2022). In addition, observational data from the Event Horizon Telescope (EHT) on the shadows of Sgr A* and M87* Akiyama et al. (2019a, b, 2022a) support the predictions of STVG theory and there are criteria to distinguish this theory from general relativity Moffat and Toth (2020). The axially symmetric solution in STVG theory provided by Moffat et al. is called the Kerr-MOG black hole Moffat (2015a). To date, several new black hole solutions related to this theory have been obtained Moffat (2021); Moffat and Toth (2021); Liu et al. (2023a). Building on the established solutions in STVG theory, extensive research has been conducted on various aspects of black holes, including their thermodynamic properties Mureika et al. (2016), geodesic motion Lee and Han (2017), shadows Moffat (2015b); Wang et al. (2019); Wei and Liu (2018); Guo et al. (2018); Sheoran et al. (2018); Zhang et al. (2024); Qiao et al. (2020); Konoplya and Zhidenko (2021); Younsi et al. (2016), and quasi-normal modes Manfredi et al. (2018); Jiang et al. (2024); Liu et al. (2023b), etc.
@@
In this work, we aim to extend the Kerr-MOG solution Moffat (2015a) to more general cases with a cosmological constant, thereby obtaining Kerr-MOG-(A)dS black holes. Given the recent groundbreaking observations by the EHT, theoretical research on the properties of black holes frequently focuses on the analysis of black hole images Chen et al. (2023); Liu et al. (2024a). Hence, we have also investigated the effects of the MOG parameter and a positive cosmological constant on black hole shadows, as well as the numerical estimation of the angular radius of the supermassive black holes Sgr A* and M87* within this theoretical framework. The remaining part of this paper is organized as follows. In Sec. II, we provide a concise overview of the STVG theory and give the Kerr-like solution of the field equations with a cosmological constant, i.e., Kerr-MOG-(A)dS black holes. In Sec. III, utilizing the geodesic equation, we derive the orbital equations for photons in the context of Kerr-MOG-dS spacetime. In Sec. IV, we examine the shadows of black holes, with a primary focus on the apparent shapes and distortions of the shadow. Also, we numerically calculate these two observables, $R_{s}$ and $\delta_{s}$, to characterize the shadow. Sec. V focuses on providing an overview of our conclusions while also contemplating the possible development and exploration of related research topics in the future.
@@ II.2 Rotating-(A)dS solutions
where $r_{m}$, $r_{h}$, and $r_{c}$ correspond to the Cauchy horizon, event horizon, and dS cosmological horizon, respectively Liu et al. (2023a, 2024b). The parameters $r_{b}$, $r_{m}$, $r_{h}$, and $r_{c}$ satisfy the condition $r_{b}+r_{m}+r_{h}+r_{c}=0$. Moreover, $r_{c}\geq r_{h}>r_{m}>0$ and $r_{b}$ is negative.
@@ III Photon orbits
where $\tau$ is an affine parameter along the null geodesics, and $\mathcal{S}$ is the Jacobi action. We can set up the Killing fields $\xi^{a}=(\partial/\partial t)^{a}$ and $\psi^{a}=(\partial/\partial\varphi)^{a}$, then the corresponding energy $\mathcal{E}$ and the angular momentum in the direction of the axis of symmetry $L$ can be defined as Wald (1984); Zeng et al. (2022); Liu et al. (2023c):
@@ IV SHADOWS OF BLACK HOLES
While AdS solutions play a crucial role in AdS/CFT correspondences, they lack empirical support from specific astronomical observations. Therefore, in this section, we primarily focus on the black hole shadows in Kerr-MOG-dS spacetime. For this, we need to adopt the analysis approach of A. Grenzebach Grenzebach et al. (2014) and the numerical ray-tracing method Wang et al. (2018a, 2021); Hu et al. (2021); Zhong et al. (2021); Chen and Jing (2024) to investigate the shadow of Kerr-MOG black holes in the presence of a positive cosmological constant.
@@ IV.1 Apparent shape
Figure 1: Shadows cast by black holes with a cosmological constant, as seen by an observer at $r_{0}=50M$ and $\vartheta_{0}=\pi/2$. The (a) and (c) are static cases $a/M=0$, corresponding to $\tilde{\Lambda}=10^{-4}$ and $\tilde{\Lambda}=10^{-3}$, respectively. The (b) and (d) are rotating case $\alpha=0.02$, corresponding to $\tilde{\Lambda}=10^{-4}$ and $\tilde{\Lambda}=10^{-3}$, respectively.
@@
Figure 2: Shadows cast by black holes with a spin $a/M=0.99$, as seen by an observer at $r_{0}=50M$ and $\vartheta_{0}=\pi/2$. The (a), (b), (c) and (d) corresponding to $\alpha=0$, $\alpha=1$, $\alpha=2$ and $\alpha=3$, respectively.
@@
Figure 3: Shadows cast by black holes with a cosmological constant, as seen by an observer at $r_{0}=50M$ and $\vartheta_{0}=\pi/2$. The (a) and (c) are spin $a/M=0.5$, corresponding to $\tilde{\Lambda}=10^{-4}$ and $\tilde{\Lambda}=10^{-3}$, respectively. The (b) and (d) are spin $a/M=0.99$, corresponding to $\tilde{\Lambda}=10^{-4}$ and $\tilde{\Lambda}=10^{-3}$, respectively.
@@
For simplicity, we consider fixing the position of the observer at $r_{0}=50M$ and $\vartheta_{0}=\pi/2$. Where the angle $\vartheta_{0}$ is defined as the inclination angle between the observer’s line of sight and the rotation axis of the black hole. From Eqs. (47) and (48), one can depict the contour of the black hole shadow under varying MOG parameter $\alpha$, spin $a$, or dimensionless cosmological constant $\tilde{\Lambda}$, which is defined as $\tilde{\Lambda}=\Lambda M^{2}$. In Fig. 1, the shadows of static black holes and rotating black holes are shown under $\tilde{\Lambda}=10^{-4}$ or $\tilde{\Lambda}=10^{-3}$. When the cosmological constant is relatively small, the shadow scale exhibits an increase with the growth of the MOG parameter. However, as the cosmological constant increases, this trend changes. This shift can be attributed to the observer’s position $r_{0}$ approaching the cosmological horizon $r_{c}$, causing the observer’s position to have a more significant effect on the shadow size than the effects of the MOG modification. In Fig. 2, we present the shadow contour of a near-extreme black hole with different $\alpha$ and $\tilde{\Lambda}$, i.e., $a/M=0.99$. The results show that, as the MOG parameter increases, the range of the photon sphere radius narrows, and its contour increasingly resembles a standard circle. When the cosmological constant is relatively small, the shadow size increases with the growth of the MOG parameter. For a fixed $a$ and $\alpha$, the apparent shape of the shadow varies with $\tilde{\Lambda}$. As $\tilde{\Lambda}$ decreases, the shadow’s size expands. In Fig. 3, the same phenomenon that is clearly visible in Fig. 1-(c) can also be observed in rotating black holes. Furthermore, when the MOG parameter increases significantly, the rotation-induced distortion of the photon orbits becomes less noticeable. It is worth noting that in Figs. 1-3, the purple solid line represents the shadow of the Schwarzschild black hole, serving as a scale reference.
@@
The contours of static versus rotating black holes will be extremely similar when the MOG parameter is sufficiently large Camilloni et al. (2024), making it difficult to infer black hole parameters solely from the black hole shadow. However, does the frame-dragging effect caused by rotation in black hole spacetimes also affect the MOG parameter? For this, we can use the numerical ray-tracing method, which adopts the same orthonormal tetrad as the analytical methods in this paper and sets the number of pixels to $n=2048$. Here, to make the manuscript concise, we would like to show our results immediately; the details of the ray-tracing method are provided in Appendix B. Figs. 4 and 5 effectively demonstrate the distortion of space by MOG black holes and the gravitational lensing effects these objects produce. Specifically, the gravitational lensing observed can be utilized to examine theories of gravity Liu et al. (2024c); Chowdhuri et al. (2023); Ghosh and Bhattacharyya (2022). We can observe that although increasing the MOG parameter makes the contours of rotating black holes approach those of static black holes, it does not affect the frame-dragging effect induced by rotation. Moreover, Figs. 4 and 5 serve as double checks for Figs. 1-(a) and 2-(b), respectively, with each displaying the same black hole contours.
@@
Figure 4: Shadows cast by static black holes ($a/M=0$) with a cosmological constant, as seen by an observer at $r_{0}=50M$ and $\vartheta_{0}=\pi/2$.
@@
Figure 5: Shadows cast by extremely rotating black holes ($a/M=0.99$) with a cosmological constant, as seen by an observer at $r_{0}=50M$ and $\vartheta_{0}=\pi/2$.
@@ IV.2 Observable measurements
Up to this point, we have determined that parameters such as spin $a$, MOG parameter $\alpha$, and dimensionless cosmological constant $\tilde{\Lambda}$ will significantly affect the apparent shape of the shadow in Kerr-MOG-dS spacetime. To characterize the shadow, it is essential to have easily measurable and reliable properties of observable measurements. We adopt two such observable measurements: the radius $R_{s}$ and distortion parameter $\delta_{s}$, following an approach analogous to the one taken in Ref. Hioki and Maeda (2009).
Figure 6: Shadows cast by black holes, as seen by an observer at $r_{0}=50M$ and $\vartheta_{0}=\pi/2$. The observable measurements for the apparent shape of the black hole are the radius of the reference circle $R_{s}$ and the distortion parameter $\delta_{s}=\mathrm{D}/R_{s}$, where $\mathrm{D}$ is the difference between the left endpoints of the reference circle and of the shadow.
The parameter $R_{s}$ characterizes the size of the shadow, corresponding to the radius of the reference circle depicted as a blue dashed line in Fig. 6. The circle passes through three points: the top position $P_{A}=(x_{t},y_{t})$, the bottom position $P_{B}=(x_{b},y_{b})$, and the point $P_{C}=(x_{r},0)$, which corresponds to the unstable retrograde circular orbit as observed from the equatorial plane by an observer. Additionally, the point $P_{O}=(x_{o},0)$ represents the center of the reference circle, and $x_{o}$ can be obtained from the coordinates of $P_{A},P_{B}$, and $P_{C}$:
@@
Figure 7: Observable measurements $R_{s}$ (left) and $\delta_{s}$ (right) as functions of $\alpha$ are presented, respectively, for $\tilde{\Lambda}=10^{-4}$ (top) and $\tilde{\Lambda}=10^{-3}$ (bottom). The observer is placed at $r_{0}=50M$ and $\vartheta_{0}=\pi/2.$
We numerically calculated these two observable measurements under the condition of fixing the observer position at $r_{0}=50M$ with an inclination angle of $\vartheta_{0}=\pi/2$. The results are presented in Fig. 7. Note that the static case is not included in the results because the contour of the shadow coincides with the reference circle, resulting in no distortion. Figs. 7-(b) and 7-(d) show that the cosmological constant $\tilde{\Lambda}$ has little effect on $\delta_{s}$. But $\delta_{s}$ decreases as the parameter $\alpha$ increases, which suggests that the Kerr-MOG-dS black hole gets less deformed compared to the Kerr-dS black hole. Figs. 7-(a) and 7-(c) clearly show the radius $R_{s}$ of the shadow. From these results, it becomes evident that when the cosmological constant $\tilde{\Lambda}=10^{-4}$, $R_{s}$ increases as the parameter $\alpha$ increases with $\alpha\in[0,3]$. However, when the cosmological constant $\tilde{\Lambda}=10^{-3}$, there is a maximum value for $R_{s}$ , corresponding to $\alpha\approx 1.815$. This is consistent with the peculiar phenomena observed in Figs. 1-(c), 3-(c), and 3-(d) describing the shadow cast by MOG-dS black holes in Sec. IV.
Figure 8: Observable measurements $R_{s}$ as functions of $\alpha$ are presented, with the positions of the observers located at $r_{0}=30M$, $r_{0}=40M$, and $r_{0}=50M$, and an observation angle of $\vartheta_{0}=\pi/2$.
To understand this phenomenon, we explored the impact of the observer’s position on $R_{s}$. For smaller cosmological constants, the domain of outer communication is immensely vast, such that the observer’s position has a negligible effect on $R_{s}$, as shown in Figs. 8-(a) and 8-(b). However, for larger cosmological constant, the domain of outer communication is more limited and further decreases with increasing MOG parameters. When the observer’s position is close to the de Sitter horizon, $R_{s}$ rapidly decreases, as shown in Fig. 8-(c) with $r=50M$ and Fig. 8-(d) with $r=40M$. The corresponding cosmological horizons are located at points from $53.744M$ to $50.378M$ with $\alpha\in[0,3]$, and from $43.686M$ to $40.214M$ with $\alpha\in[0,3]$, respectively.
For a approximatively estimation, utilizing the Kerr-MOG-dS metric (22), we calculate the angular radius of a black hole shadow, defined as $\theta_{\text{BH}}=R_{s}\frac{\mathcal{M}}{D_{O}}$, with $D_{O}$ representing the distance from the observer to the black hole. Specifically, for a black hole possessing a mass of $\mathcal{M}$ and situated $D_{O}$ away from the observer, we express the angular radius $\theta_{\text{BH}}$ quantitatively as $\theta_{\text{BH}}=9.87098\times 10^{-6}R_{s}\left(\frac{\mathcal{M}}{M_{\odot}}\right)\left(\frac{1\text{kpc}}{D_{O}}\right)\mu\text{as}$ Amarilla and Eiroa (2012).
@@
$\tilde{\Lambda}=0.0000$ 24.737 26.747 28.738 30.712 32.670
$\tilde{\Lambda}=0.0001$ 23.644 25.562 27.460 29.342 31.208
$\tilde{\Lambda}=0.0002$ 22.497 24.317 26.118 27.902 29.671
$\tilde{\Lambda}=0.0003$ 21.288 23.004 24.701 26.382 28.047
$\tilde{\Lambda}=0.0000$ 24.751 26.760 28.750 30.722 32.679
$\tilde{\Lambda}=0.0001$ 23.658 25.573 27.470 29.351 31.215
$\tilde{\Lambda}=0.0002$ 22.510 24.328 26.127 27.909 29.677
$\tilde{\Lambda}=0.0003$ 21.298 23.013 24.709 26.388 28.053
@@
$\tilde{\Lambda}=0.0000$ 19.859 21.473 23.081 24.656 26.229
$\tilde{\Lambda}=0.0001$ 18.982 20.522 22.046 23.556 25.055
$\tilde{\Lambda}=0.0002$ 18.062 19.522 20.968 22.401 23.821
$\tilde{\Lambda}=0.0003$ 17.090 18.457 19.816 21.164 22.517
$\tilde{\Lambda}=0.0000$ 19.871 21.484 23.064 24.664 26.236
$\tilde{\Lambda}=0.0001$ 18.993 20.531 22.054 23.563 25.061
$\tilde{\Lambda}=0.0002$ 18.071 19.531 20.975 22.406 23.826
$\tilde{\Lambda}=0.0003$ 17.099 18.475 19.837 21.185 22.521